OceanGate sells external hard drives for $261 each. Its total fixed costs are $46 million, and its variable costs per unit are $170. The corporate tax rate is 35%. If the economy is strong, the firm will sell 6 million drives, but if there is a recession, it will sell only half as many. a. What will be the percentage decline in sales if the economy enters a recession? (Negative value should be indicated by a minus sign. Round your answer to 2 decimal places.) Percentage decline in sales b. What will be the percentage decline in profits if the economy enters a recession? (Negative value should be indicated by a minus sign. Round your answer to 2 decimal places.) Percentage decline in profits c. Calculate the operating leverage of this firm? (Round your answer to 2 decimal places.) DOL
Answer #1

Answer (a):

Based on information given, sales, variable cost, contribution, net profit are calculated below for both the scenarios:

Total amount (in million $ Per Unit for 6 million units for 3 million units Sale Variable cost Contribution Fixed cost Profit

Percentage decline in sales:

In terms of units (in million) = (3 - 6) / 6 = - 50.00%

In terms of dollar sales value (in million $) = (783 - 1566) / 1566 = - 50.00%

Percentage decline in sales )% 0.00

Answer (b):

Percentage decline in profit (in million $) = (147.55 - 325) / 325 = -54.60%

Percentage decline in profit ! S4.601%

Answer (c):

Operating Leverage = (Sales - Variable cost) / (Sales - Variable cost - Fixed cost)

= (1566 - 1020) / (1566 -1020 - 46) = 1.092 = 1.09

Or

Operating Leverage = % Change in EBIT / % Change in Sales

% Change in sales (as calculated in answer 1 above) = -50.00%

% Change in EBIT = (227 - 500) / 500 = - 54.60%

Operating Leverage = - 54.60% / -50.00% = 1.092 = 1.09

DOL 1.09